FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF ARGON ACETONE

ALISON R. KEIMOWITZ, MICHAELEEN R. MUNROW, and \underline{STEWART E. NOVICK}, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The microwave spectra of the argon acetone van der Waals complex has been obtained in a pulsed jet Fourier transform microwave spectrometer. The geometry of the complex and the effect of van der Waals bonding on the internal rotation of the two methyl groups will be discussed. In addition, the order of the Stark effect in the various internal rotation states of the complex and the acetone monomer will be addressed.